If zis not 0 prove that (2)-1 = 2--1
Answer #1

let z=nrig ... : them, z = n-ly (?)*= {-iy)? وأ- e pylarig). ntin saty? Now. I = Inting - here in) = ntin het nga There que,Alternative let z = r (Cerotismo), nato t = ro (Sno-i simo) que fomoio -i Simon = pant Cap-i süno) = pol (como + is wo). Agai